# Towards Embodied Agentic AI: Review and Classification of LLM- and VLM-Driven Robot Autonomy and Interaction

### v2

@@ Abstract
Foundation models, including large language models (LLMs) and vision-language models (VLMs), have recently enabled novel approaches to robot autonomy and human-robot interfaces. In parallel, vision-language-action models (VLAs) or large behavior models (LBMs) are increasing the dexterity and capabilities of robotic systems. This survey paper focuses on those works advancing towards agentic applications and architectures. This includes initial efforts exploring GPT-style interfaces to tooling, as well as more complex system where AI agents are coordinators, planners, perception actors, or generalist interfaces. Such agentic architectures allow robots to reason over natural language instructions, invoke APIs, plan task sequences, or assist in operations and diagnostics. In addition to peer-reviewed research, due to the fast-evolving nature of the field, we highlight and include community-driven projects, ROS packages, and industrial frameworks that show emerging trends. We propose a taxonomy for classifying model integration approaches and present a comparative analysis of the role that agents play in different solutions in today’s literature.
{strip}
Figure 1: Illustration of different approaches to the integration of LLMs, VLMs and other foundation models into robotic systems. This survey focuses on the integration perspective and the role of such models within the wider robotic system where it is deployed. Our goal is to discuss the current state-of-the-art within the context of Embodied Agentic AI applications.
@@ I Introduction
Several surveys have addressed the broader intersection of foundation models and robotics [1, 2, 3, 4, 5], with a focus on multimodal architectures and their components and primarily emphasizing the general use of multimodal models in robotic decision-making and high-level planning, as well as end-to-end learning frameworks for low-level control in domain-specific settings such as manipulation. However, with the very recent rise of emergence of embodied and generalist AI agents, no prior survey to date has examined the emerging design patterns in which AI agents interface with existing control software, libraries, or middleware. Moreover, many practical systems (GitHub-hosted projects, ROS packages, or startup prototypes) remain underrepresented in the literature despite their increasing real-world relevance and impact.
@@
  • We propose a taxonomy across two dimensions: integration approach to foundation models (Section III, see also Fig. 1), and agent role (Section IV, see also Fig. 3).
@@ II Towards Embodied Agentic AI
The concept of an embodied robotic agent was introduced in the late 1980s, grounded in the idea that intelligent behavior emerges from a system’s physical embodiment and its continuous sensorimotor interaction with the environment [17]. This laid the foundation for early end-to-end neural models that directly mapped perception to action, enabling autonomous control without explicit symbolic reasoning [18, 19, 20]. The emergence of LLMs has significantly expanded the functional scope of autonomous agents, shifting them from narrowly scoped, rule-driven systems to more flexible entities capable of generalization, contextual reasoning, and adaptive behavior [21]. At pre-ChatGPT era, Code as Policies (CaP) [6] introduced robot-executable language model generated programs (LMPs), enabling reactive and vision-based control via indirect tool calling through code generation.
@@
Early exploration in this domain focused on natural language interfaces for ROS 2, including ros2ai [22], an extension to augment the ROS 2 command line interface with LLM, and ROScribe [23], a tool using LLMs to generate ROS codebases from a high-level description (e.g. a user describe desired ROS graph (sensors, topics, and functionality) and ROScribe will output Python code for ROS nodes and launch files to implement that spec). Furthermore, The ROS-LLM framework [24] was proposed to enable embodied intelligence applications. It provides a template to define your robot’s API in a config file, then handles the prompt engineering and calls to the OpenAI API, effectively acting like a customizable agent. Concurrently, Microsoft’s ChatGPT for Robotics [7] introduced a modular approach to LLM-based robot control and programming, defined a high-level API specific to each robot, leveraging prompt engineering and pre-defined function libraries to enable natural language control across diverse tasks and simulation environments. During this phase, most implementations adopted a chatbot-style design, deploying LLMs to interpret free-form textual inputs and generate appropriate commands for general robot integration.
@@
Building on these advances, 2024 marked a transformative shift with the development of specialized LLM interfaces for robotics and VLA systems, unlocked unprecedented generalization and semantic reasoning capabilities in robotic control systems. These frameworks enabled agents to process open-ended natural language instructions, ground them in visual perception, and execute corresponding actions [14]. A notable breakthrough was RT-2 [8], which introduced a novel VLA framework by co-fine-tuning pretrained VLMs on both robotic trajectory data and internet-scale vision-language tasks. $\pi$[12] advanced the field by introducing real-time, continuous control via flow-matched diffusion policies, unifying perception, reasoning, and motion generating in a single differentiable framework.
@@
Simultaneously, agentic middleware frameworks such as ROSA (Robot Operating System Agent) [9], RAI (Robotic AI Agent) [10], and BUMBLE [11] emerged, bridging the gap between LLM-based reasoning and traditional robotic middleware, in the pursuit of embodied AI. Initially open-sourced in late 2023, ROSA represents a significant advancement in AI-driven robotic control by implementing an LLM agent built on the LangChain framework and the ReAct (Reasoning and Acting) agent paradigm. Its core innovation was the abstraction of ROS operations into tool-enabled Python functions - including listing available nodes, reading sensor data, moving a joint, etc. This allowed ROSA to translate natural language commands into validated robot actions. ROSA also embedded safety mechanisms like parameter validation, constraint enforcement, and optional human approval for critical actions. The authors demonstrate ROSA across heterogeneous platforms, including JPL’s NeBula-Spot quadruped and NVIDIA Isaac Sim environments. However, its architecture remains tightly coupled to ROS-based middleware, limiting interoperability with non-ROS systems and constraining deployment in broader contexts. RAI is another approach in this field as a flexible embodied multi-agent framework designed to integrate LLM reasoning with robotic systems (e.g. ROS 2). Its architecture introduces a novel distributed paradigm where specialized LLM agents, including perception modules, task planners, motion controllers, and safety monitors, collaborate through well-defined roles to enable concurrent, real-time task execution while maintaining system safety and reliability. The framework’s core components (Agents, Connectors, and Tools) provide integrated capabilies for multimodel sensing and actuation, retrieval‑augmented generation through vector stores and pre‑configured agent types (voice, conversational, state‑based). The framework demonstrates particular advancement in dynamic environments, featuring capabilities like online replanning and failure recovery. Proven on both physical (Husarion ROSBot XL) and simulated (tractor and manipulator) platforms. However, the framework is constrained by the limited spatial reasoning and unreliable self-correction of LLM-based agents, which repeatedly led to failures in understanding physical boundaries, object interactions, and task completion. BUMBLE (BUilding-wide MoBiLE Manipulation) is another recent contribution that uses a unified VLM-based framework for building-wide mobile manipulation. It integrates open-world perception, dual-layered memory systems, and a wide range of motor skills, allowing effective operation over extended spatial and temporal horizons. Unlike frameworks such as ROSA, which rely heavily on ROS-based middleware integration, BUMBLE features tighter coupling with navigation systems through the use of predefined landmark images. While this provides efficient and precise navigation, it also introduces limitations, such as requiring prior manual collection of landmark imagery. Nevertheless, BUMBLE’s combination of perception and action capabilities presents a significant advance in addressing the complexity of real-world environment.
@@
Another notable development is the concept of an MCP (Model Context Protocol) server, driving agentic use and quick integration of new capabilities into existing tools (e.g., Anthropic’s Claude or AI-driven code editors like Cursor). Their penetration in robotics is so far community-driven, with several projects openly available on GitHub. The ROS-MCP library [25], for example, connects to AI assistants such as Claude and enables interaction with installed ROS 2 applications. The capabilities include (i) launching UI applications, including the Gazebo simulator and the RViz visualizer; (ii) ROS resource management (topics, nodes); (iii) ROS system interaction (publishing to topics, calling services, send action goals); (iv) environment debugging; and (v) process management (e.g., cleaning up running ROS 2 processes). This represents a paradigm shift from other frameworks that provide a complete application system such as ROSA and RAI, with MCP servers taking a plugin-based approach instead, requiring standard AI assistant applications. Another example is the ros-mcp-server [25], which uses rosbridge to connect to either ROS 1 or ROS 2 systems through WebSockets, and contains specific interfaces to odometry and velocity control topics. A benefit of leveraging assistants like Claude is that they already have other built-in tooling, such as web search, and the ability to run code internally in a sandboxed environment. This opens the possibility to more advanced behaviors without additional implementation.
@@
A final and more recent open-source framework for agentic AI in robotics is driven by the startup OpenMind [16]. OpenMind’s OM1 introduces a modular, hardware-agnostic AI runtime intended for various robotic platforms. Compared to earlier frameworks such as ROSA and RAI, which primarily target integration with ROS-based middleware, OM1 differentiates itself with its decentralized FABRIC coordination protocol. FABRIC enables secure identity management and interoperability among heterogeneous robotic systems, diverging from the centralized orchestration common in earlier systems. While OM1 utilizes Python-based configurations and incorporates perception-action loops similar to VLA-driven systems like RT-2, it notably emphasizes decentralization. Overall, OM1 represents a step toward more collaborative and flexible robotic ecosystems.
@@
Figure 2 illustrates in a timeline format some of the key milestones and most representative works of the past three years. The top part of the figure focuses on agentic-oriented approaches, while the bottom part lists other relevant models that have shaped the advances of the field. A more comprehensive list of works, both academic papers and community projects, is available in Table II.
TABLE I: Selection of most representative works spanning the time since the release of ChatGPT and the different integration approaches proposed in this survey.
@@
CaP [6] Robot-Centric formulation of language model programs 2022 Protocol Indirect tool through coding Variable logging LLM N.A
ros2ai [22] ROS 2 command line interface extension with LLM 2023 Protocol ROS 2 CLI LLM N.A
ROS-LLM [24] ROS framework for embodied intelligence applications 2023 Protocol ROS actions/services or code API Logging LLM N.A
ROScribe [23] Create ROS packages using LLMs 2023 Protocol/ Interface Custom tools - LLM - N.A
ChatGPT Robotics [7] High-level functions library for robotics application 2023 Protocol/ Interface Custom tools LLM Sys. prompt & tools
ROSA [9] ROS Agent, interact with ROS-based robotics systems 2024 Interface Custom tools Chat history VLM Sys. prompt & tools
RAI [10] AI agent framework for embodied AI 2024 Interface/ Framework Custom toolsets - VLM Custom tools
Embodied Agents [26] ROS 2 based framework for robot to understand models 2024 Interface Custom tools Semantic Memory VLM Sys. prompt & tools
BUMBLE [11] Framework for open-world perception system 2024 Interface Predefined skill library VLM Custom skills
AutoRT [27] Models for Orchestration of Robotic Agents 2024 Orchestration Pretrained RT policies - VLM/ VLA Pretrained policies
ROS-MCP [28, 25] Converts natural language to ROS command 2025 Protocol Custom tools - LLM - Custom tools
Gemini Robotics [13] Generalist VLA model 2025 Embedded Tool library - VLA - -
MoManip VLA [29] VLA for general mobile manipulation 2025 Embedded Pretrained policies - VLA ✓soon Pretrained policies
ARNA [30] General-Purpose Robotic Navigation via LVLM 2025 Embedded Custom toolsets Multimodal memory VLM ✓soon Custom tools
OpenMind OM1 [16] Modular AI runtime for agents and robots 2025 Interface/ Framework Custom toolsets Chat history VLM Custom tools
UROSA [31] Distributed AI Agents for Underwater Robot Autonomy 2025 Framework Custom toolsets/ On the fly code generation Logging VLM/ LLM Sys. prompts
TABLE II: Selection of agentic AI frameworks relevant to this survey but not developed for or tested on robotic systems.
Voyager [32] LLM-powered embodied lifelong learning agent 2023 Embedded Skill library + iterative prompt LLM - -
Odyssey [33] Agents with Open-World Skills 2025 Embedded Predefined skill library LLM - Predefined skill

### v3

@@ Abstract
Foundation models, including large language models (LLMs) and vision-language models (VLMs), have recently enabled novel approaches to robot autonomy and human-robot interfaces. In parallel, vision-language-action models (VLAs) or large behavior models (LBMs) are increasing the dexterity and capabilities of robotic systems. This survey paper reviews works that advance agentic applications and architectures, including initial efforts with GPT-style interfaces and more complex systems where AI agents function as coordinators, planners, perception actors, or generalist interfaces. Such agentic architectures allow robots to reason over natural language instructions, invoke APIs, plan task sequences, or assist in operations and diagnostics. In addition to peer-reviewed research, due to the fast-evolving nature of the field, we highlight and include community-driven projects, ROS packages, and industrial frameworks that show emerging trends. We propose a taxonomy for classifying model integration approaches and present a comparative analysis of the role that agents play in different solutions in today’s literature.
@@ I Introduction
Several surveys have addressed the broader intersection of foundation models and robotics [firoozi2025foundation, urain2024deep, zeng2023large, xiao2025robot, kim2024survey], with a focus on multimodal architectures and their components and primarily emphasizing the general use of multimodal models in robotic decision-making and high-level planning, as well as end-to-end learning frameworks for low-level control in domain-specific settings such as manipulation. However, with the very recent rise of emergence of embodied and generalist AI agents, no prior survey to date has examined the emerging design patterns in which AI agents interface with existing control software, libraries, or middleware. Moreover, many practical systems (GitHub-hosted projects, ROS packages, or startup prototypes) remain underrepresented in the literature despite their increasing real-world relevance and impact.
@@
  • We propose a taxonomy across two dimensions: integration approach to foundation models (Section III, see also Fig. LABEL:fig:concept), and agent role (Section IV, see also Fig. 3).
  • We include the section (§V) presenting a practical design toolkit to guide implementation and prototyping.
  • Finally, the paper concludes with a detailed discussion (§VI) that highlights key open questions and outlines promising directions for future research in agentic embodied AI.
@@ II Towards Embodied Agentic AI
The concept of an embodied robotic agent was introduced in the late 1980s, grounded in the idea that intelligent behavior emerges from a system’s physical embodiment and its continuous sensorimotor interaction with the environment [brooks2003robust]. This laid the foundation for early end-to-end neural models that directly mapped perception to action, enabling autonomous control without explicit symbolic reasoning [bojarski2016end, sermanet2018time, laskin2020reinforcement]. The emergence of LLMs has significantly expanded the functional scope of autonomous agents, shifting them from narrowly scoped, rule-driven systems to more flexible entities capable of generalization, contextual reasoning, and adaptive behavior [yao2023react]. At pre-ChatGPT era, Code as Policies (CaP) [liang2022code] introduced robot-executable language model generated programs (LMPs), enabling reactive and vision-based control via indirect tool calling through code generation.
@@
Early exploration in this domain focused on natural language interfaces for ROS 2, including ros2ai [ros2ai], an extension to augment the ROS 2 command line interface with LLM, and ROScribe [roscribe], a tool using LLMs to generate ROS codebases from a high-level description (e.g. a user describes a desired ROS graph (sensors, topics, and functionality) and ROScribe will output Python code for ROS nodes and launch files to implement that spec). Furthermore, the ROS-LLM framework [rosllm] was proposed to enable embodied intelligence applications. It provides a template to define your robot’s API in a config file, then handles the prompt engineering and calls to the OpenAI API, effectively acting like a customizable agent. Concurrently, Microsoft’s ChatGPT for Robotics [vemprala2024chatgpt] introduced a modular approach to LLM-based robot control and programming, defined a high-level API specific to each robot, leveraging prompt engineering and pre-defined function libraries to enable natural language control across diverse tasks and simulation environments. During this phase, most implementations adopted a chatbot-style design, deploying LLMs to interpret free-form textual inputs and generate appropriate commands for general robot integration.
@@
Building on these advances, 2024 marked a transformative shift with the development of specialized LLM interfaces for robotics and VLA systems, unlocked unprecedented generalization and semantic reasoning capabilities in robotic control systems. These frameworks enabled agents to process open-ended natural language instructions, ground them in visual perception, and execute corresponding actions [bjorck2025gr00t]. A notable breakthrough was RT-2 [zitkovich2023rt], which introduced a novel VLA framework by co-fine-tuning pretrained VLMs on both robotic trajectory data and internet-scale vision-language tasks. $\pi$[black2410pi0] advanced the field by introducing real-time, continuous control via flow-matched diffusion policies, unifying perception, reasoning, and motion generating in a single differentiable framework.
@@
Simultaneously, agentic middleware frameworks such as ROSA (Robot Operating System Agent) [royce2025enabling], RAI (Robotic AI Agent) [rachwal2025rai], and BUMBLE [shah2024bumble] emerged, bridging the gap between LLM-based reasoning and traditional robotic middleware, in the pursuit of embodied AI. Initially open-sourced in late 2023, ROSA represents a significant advancement in AI-driven robotic control by implementing an LLM agent built on the LangChain framework and the ReAct (Reasoning and Acting) agent paradigm. Its core innovation was the abstraction of ROS operations into tool-enabled Python functions - including listing available nodes, reading sensor data, moving a joint, etc. This allowed ROSA to translate natural language commands into validated robot actions. ROSA also embedded safety mechanisms like parameter validation, constraint enforcement, and optional human approval for critical actions. The authors demonstrate ROSA across heterogeneous platforms, including JPL’s NeBula-Spot quadruped and NVIDIA Isaac Sim environments. However, its architecture remains tightly coupled to ROS-based middleware, limiting interoperability with non-ROS systems and constraining deployment in broader contexts. RAI is another approach in this field as a flexible embodied multi-agent framework designed to integrate LLM reasoning with robotic systems (e.g. ROS 2). Its architecture introduces a novel distributed paradigm where specialized LLM agents, including perception modules, task planners, motion controllers, and safety monitors, collaborate through well-defined roles to enable concurrent, real-time task execution while maintaining system safety and reliability. The framework’s core components (Agents, Connectors, and Tools) provide integrated capabilies for multimodel sensing and actuation, retrieval‑augmented generation through vector stores and pre‑configured agent types (voice, conversational, state‑based). The framework demonstrates particular advancement in dynamic environments, featuring capabilities like online replanning and failure recovery. Proven on both physical (Husarion ROSBot XL) and simulated (tractor and manipulator) platforms. However, the framework is constrained by the limited spatial reasoning and unreliable self-correction of LLM-based agents, which repeatedly led to failures in understanding physical boundaries, object interactions, and task completion. BUMBLE (BUilding-wide MoBiLE Manipulation) is another recent contribution that uses a unified VLM-based framework for building-wide mobile manipulation. It integrates open-world perception, dual-layered memory systems, and a wide range of motor skills, allowing effective operation over extended spatial and temporal horizons. Unlike frameworks such as ROSA, which rely heavily on ROS-based middleware integration, BUMBLE features tighter coupling with navigation systems through the use of predefined landmark images. While this provides efficient and precise navigation, it also introduces limitations, such as requiring prior manual collection of landmark imagery. Nevertheless, BUMBLE’s combination of perception and action capabilities presents a significant advance in addressing the complexity of real-world environment.
@@
Another notable development is the concept of an MCP (Model Context Protocol) server, driving agentic use and quick integration of new capabilities into existing tools (e.g., Anthropic’s Claude or AI-driven code editors like Cursor). Their penetration in robotics is so far community-driven, with several projects openly available on GitHub. The ROS-MCP library [rosmcpserver], for example, connects to AI assistants such as Claude and enables interaction with installed ROS 2 applications. The capabilities include (i) launching UI applications, including the Gazebo simulator and the RViz visualizer; (ii) ROS resource management (topics, nodes); (iii) ROS system interaction (publishing to topics, calling services, send action goals); (iv) environment debugging; and (v) process management (e.g., cleaning up running ROS 2 processes). This represents a paradigm shift from other frameworks that provide a complete application system such as ROSA and RAI, with MCP servers taking a plugin-based approach instead, requiring standard AI assistant applications. Another example is the ros-mcp-server [rosmcpserver], which uses rosbridge to connect to either ROS 1 or ROS 2 systems through WebSockets, and contains specific interfaces to odometry and velocity control topics. A benefit of leveraging assistants like Claude is that they already have other built-in tooling, such as web search, and the ability to run code internally in a sandboxed environment. This opens the possibility to more advanced behaviors without additional implementation. Additional implementation developed by robotmcp [robotmcp] provides a unified interface that connects LLMs with existing ROS/ROS 2 systems without requiring source code changing, enabling true two-way integration where natural language commands are translated into ROS actions while AI can simultaneously monitor the robot states and sensor data in real time.
@@
A final and more recent open-source framework for agentic AI in robotics is driven by the startup OpenMind [openmind]. OpenMind’s OM1 introduces a modular, hardware-agnostic AI runtime intended for various robotic platforms. Compared to earlier frameworks such as ROSA and RAI, which primarily target integration with ROS-based middleware, OM1 differentiates itself with its decentralized FABRIC coordination protocol. FABRIC enables secure identity management and interoperability among heterogeneous robotic systems, diverging from the centralized orchestration common in earlier systems. While OM1 utilizes Python-based configurations and incorporates perception-action loops similar to VLA-driven systems like RT-2, it notably emphasizes decentralization. Overall, OM1 represents a step toward more collaborative and flexible robotic ecosystems.
@@
Figure 2 illustrates in a timeline format some of the key milestones and most representative works of the past three years. The top part of the figure focuses on agentic-oriented approaches, while the bottom part lists other relevant models that have shaped the advances of the field. A more comprehensive list of works, both academic papers and community projects, is available in Table II. The categorization of integration approaches is discussed in detail in Section III. The toolset/skillset primarily reflects the mechanisms through which each system executes actions. In most works, these actions are realized through custom toolsets, referring to predefined tool classes within agentic frameworks or task-specific functions designed by developers. In other cases—particularly for vision-based systems—the actions are driven by pretrained models or policy networks that interpret sensory inputs to perform manipulation or navigation tasks. The memory component, a fundamental aspect of agentic systems, is generally implemented in relatively common forms such as chat history or logging-based storage. In a few recent works, memory modules are extended to structured databases or multimodal memory architectures that store both textual and visual context. The world model refers not only to the representation of the physical environment, objects, and robot dynamics but also to high-level behavioral and reasoning constructs, such as system prompts that define goals, constraints, and operational rules. While the concept of a world model is closely related to the notion of adaptation, as both contribute to an agent’s ability to generalize and reason about novel situations, the world model primarily supports environmental understanding, whereas adaptability reflects the system’s capacity to update or extend its skills in response to new tasks.
TABLE I: Comparison of Integration Approaches
Aspect Protocol-Focused Integration Interface / Agentic Integration Orchestration-Oriented Integration
Main role LLM maps user text directly to tool/API calls LLM acts as user-facing agent that reasons and executes actions LLM (or small crew) coordinates multiple agents, skills, or robots
Tooling focus Individual API or CLI calls, simple wrappers MCP or internal tools, loops or concurrent tool calls LangChain/LangGraph, or custom multi-agent orchestration pipelines
Typical frameworks MCP servers, ROS CLI bindings LangChain, LangGraph (ReAct agents), ROSA, RAI LangGraph, custom orchestrators (e.g., AutoRT)
User interaction Direct query–response Conversational, human-in-the-loop reasoning Limited direct dialogue; user gives high-level goal, orchestrator manages subtasks internally
Physical interaction Minimal or indirect (via one API) Yes: calls tools that affect sensors or actuators Indirect: delegates real-world execution to sub-agents or robots
Agent focus Translation between natural language and system commands Reasoning + execution Task assignment, scheduling, and monitoring
Analogy Command translator Assistant / operator Project manager / mission control
TABLE II: Selection of most representative works spanning the time since the release of ChatGPT and the different integration approaches proposed in this survey.
@@
CaP [liang2022code] Robot-Centric formulation of language model programs 2022 Protocol Indirect tool through coding Variable logging LLM N.A
ros2ai [ros2ai] ROS 2 command line interface extension with LLM 2023 Protocol ROS 2 CLI LLM N.A
ROS-LLM [rosllm] ROS framework for embodied intelligence applications 2023 Protocol ROS actions/services or code API Logging LLM N.A
ROScribe [roscribe] Create ROS packages using LLMs 2023 Protocol/ Interface Custom tools - LLM - N.A
ChatGPT Robotics [vemprala2024chatgpt] High-level functions library for robotics application 2023 Protocol/ Interface Custom tools LLM Sys. prompt & tools
ROSA [royce2025enabling] ROS Agent, interact with ROS-based robotics systems 2024 Interface Custom tools Chat history VLM Sys. prompt & tools
RAI [rachwal2025rai] AI agent framework for embodied AI 2024 Interface/ Framework Custom toolsets - VLM Custom tools
Embodied Agents [embodiedagent] ROS 2 based framework for robot to understand models 2024 Interface Custom tools Semantic Memory VLM Sys. prompt & tools
BUMBLE [shah2024bumble] Framework for open-world perception system 2024 Interface Predefined skill library Multimodal memory VLM Custom skills
AutoRT [ahn2024autort] Models for Orchestration of Robotic Agents 2024 Orchestration Pretrained RT policies - VLM/ VLA Pretrained policies
ROS-MCP [rosmcp, rosmcpserver, robotmcp] Converts natural language to ROS command 2025 Protocol Custom tools - LLM - Custom tools
Gemini Robotics [team2025gemini] Generalist VLA model 2025 Embedded Tool library - VLA - -
MoManip VLA [wu2025momanipvla] VLA for general mobile manipulation 2025 Embedded Pretrained policies - VLA ✓soon Pretrained policies
ARNA [lange2025general] General-Purpose Robotic Navigation via LVLM 2025 Embedded Custom toolsets Multimodal memory VLM ✓soon Custom tools
OpenMind OM1 [openmind] Modular AI runtime for agents and robots 2025 Interface/ Framework Custom toolsets Chat history VLM Custom tools
explainable ros [explainable] Explanations of robot behavior using RAG method 2024 Protocol Custom toolsets Multimodal vector database LLM/ VLM N.A
LLama ros [llama_ros] Integrate llama.cpp with ROS 2 to enable LLM inference 2024 Protocol/ Interface Custom toolsets LLM/ VLM Custom tools
Stretch AI [stretch_ai] Orchestrates skills for language-directed mobile manipulation 2024 Orchestration Custom Policies for manipulation Dynamic memory LLM Custom Skillsets
UROSA [urosa] Distributed AI Agents for Underwater Robot Autonomy 2025 Framework Custom toolsets/ On the fly code generation Logging VLM/ LLM Sys. prompts